# Free-style and Fast 3D Portrait Synthesis

### v1

@@ 4.4 Comparison experiments
Figure 6: Comparison results between our method and baselines. The text prompts and Ip2p editing results for each style are displayed on the top.
Baselines. We divide the baselines into three categories. 1) Image style transfer + GAN Inversion. 3DGAN Inversion [9] is selected as the baseline, which inverts the stylized 2D portrait. 2) Text-to-3D. DreamFusion [^1] https://github.com/ashawkey/stable-dreamfusion [41] is a representative method that generates 3D images based on the text prompts. 3) 3D GAN + 2D portrait stylization. We directly set our method without I2T as the baseline.
Table 1: Comparison of running time.
method average time
3DGAN Inversion [9] $\sim$ 2 mins
DreamFusion [41] $\sim$ 40 mins
Ours $\sim$ 3 mins
The results of our method compared with baselines are shown in Fig. 6. The portraits with large stylistic variations cannot be inverted well using the 3DGAN Inversion method, because the 3D portraits synthesized by such method cannot escape from the domain of the pre-trained 3DGAN model. DreamFusion, the representative of the Text-to-3D method, is able to optimize the model to synthesize 3D portraits based on text prompts, such as "a Elf", but the quality of generated results is low, while for more specific portrait styles, such as "sand painting style", no results can be synthesized. Besides, we display our results without I2T module, i.e., samples in the few-shot dataset. It can be seen there are variations in the synthesized portraits in each view, and 3D consistency cannot be guaranteed. In contrast, our method is able to synthesize high quality 3D portraits that satisfy both 3D consistency and stylization. In addition, our method also has a significant advantage in running time, as shown in Tab. 1. Our method is able to fine-tune a single-style model in about 3 minutes, while DreamFusion needs about 40 minutes to train on a single text. Although 3DGAN Inversion costs comparable running time to ours, its portrait generation quality is poor and fails to generate free-style portraits.
@@ 4.5 Ablation study
Table 2: Quantitative results on I2T network.
@@
Ours 0.306
We conduct the ablation studies of our method from three aspects, which are shown in Fig. 7. First we perform the ablation of the I2T network in (a) of Fig. 7. 1) When we remove the I2T network and instead add two modulation layers directly to the original EG3D triplane generator (only the modulation layers are trainable, original EG3D triplane generator is fixed), we can achieve some stylization effect. However, the generation quality is limited by the lightweight modulation layer, especially applied to multi-style and multi-identity synthesis. Furthermore, we perform a quantitative evaluation, as shown in Tab. 2. We use the CLIP [42] model to calculate the feature cosine similarity between the portrait image and the text description (e.g. "a elf face") to measure text-matching degree. Our method achieves the best CLIP score. 2) When using our I2T network without pre-training, training I2T from scratch on single-style dataset will cause mode collapse. Because it is difficult to establish the mapping from the image to the triplane representation only using 3D inconsistent data. 3) When the $\mathbf{w}_{s}$ code of our I2T network is a vector without spatial dimension, the learned 3D portrait is blurred and lacks details.
Figure 7: The ablation studies of our approach. (a) 3D portrait synthesis results in I2T network ablation experiment. (b) 3D shape outputs in dr loss ablation experiment. (c) Ip2p generated results in inference optimization ablation experiment.
We ablate the density regularization (dr) loss used in model training, shown in (b) of Fig. 7. When the density regularization is not used, the synthesized portrait shape will have rough surfaces. In addition, when we do not optimize the Ip2p inference, some basic attribute editing prompts generate poor results, such as "The man with long hair", and the stylization varies greatly at different views for some prompts, such as "a wizard" (see (c) in Fig. 7).
@@
5 Limitation
Our approach is based on two powerful pre-trained generative priors, which are the basis of our method’s ability to synthesize high-quality stylized 3D portraits. At the same time, our method is limited by both priors, especially Ip2p, which is unable to achieve perfect 3D-consistent portrait stylization in different viewpoints, so the final synthesized 3D portrait of our method differs slightly from the style generated by Ip2p. Meanwhile, some stylistic changes that differ significantly from the human portrait shape, such as "Iron Man" and "Stormtrooper" in Fig. 8, result in poorer quality 3D portraits compared to human portraits. Because when Ip2p generates this type of style, the stylization effect will be more different under different views.
@@ 6 Conclusions
This paper proposes a novel free-style and fast 3D portrait synthesis framework. Our method is based on two pre-trained generative priors, EG3D and Ip2p. We optimize Ip2p inference in order to stylize portraits at different views more freely and stable. We can quickly construct a few-shot training set of stylized portraits using the two generative priors, and fast fine-tune the EG3D prior. Moreover, We replace the original triplane generator in EG3D with a trainable I2T network to help fine-tune EG3D more efficiently. We also build a multi-style and multi-identity portrait dataset to demonstrate the scalability of our approach. A large number of high-quality 3D portrait synthesis results and comparison experiments with baselines show the superiority of our method.

### v3

@@ 4.3 Comparison experiments
Figure 7: Qualitative comparison results between our method and baselines. Reference image represents the result after Ip2p style transformation, which is used as the input of the 3D GAN Inversion methods. The part with no result means that the method is unable to generate a image.
Table 1: User Study, quantitative evaluation and running time of different methods.
method Text-image Similarity Image Quality 3D Consistency CLIP Score Average Time
DreamFusion [34] 2.82 2.24 3.18 0.304 $\sim$ 40 mins
3DGAN-Inv [23] 2.04 1.90 2.59 0.229 $\sim$ 4 mins
HFGI3D [42] 3.40 3.54 3.67 0.310 $\sim$ 10 mins
Ours 4.05 4.27 4.34 0.332 $\sim$ 3 mins
Baselines. We divide the baselines into two categories. 1) Text-to-3D. DreamFusion [^1] https://github.com/ashawkey/stable-dreamfusion [34] is a representative method that generates 3D images based on the text prompts. 2) Image style transfer + 3D GAN Inversion. 3DGAN-Inv [23] and HFGI3D [42] are two SOTA methods of 3D GAN inversion. We use them with Instruct-pix2pix as the baselines.
The results of our method compared with baselines are shown in Fig. 7. DreamFusion, the representative of the Text-to-3D method, is able to optimize the model to synthesize 3D portraits based on text prompts, such as ”A Disney style Elf”, but the generated results have low quality, while for more specific portrait styles, such as ”sand painting style”, no results can be synthesized. What’s more, the portraits with large stylistic variations cannot be inverted well using the 3DGAN-Inv, because the 3D portraits synthesized by this method cannot escape from the domain of the pre-trained 3D-aware GAN model. The results of HFGI3D are better, but the 3D shape is destroyed after inversion, as shown in the Elf’s face. At the same time, for some more difficult samples, this method cannot produce effective results, such as the last two lines. In contrast, our method is able to synthesize high quality 3D portraits that satisfy both 3D consistency and stylization.
For quantitative evaluation, we conduct a user study. We invite 30 volunteers to evaluate each method from three perspectives, namely the text-image similarity, the quality of the generated images, and the 3D consistency of the results. Each item is scored on a 5-point scale, and the average is calculated as the final result. As shown in Tab. 1, our approach achieves the best scores under each dimension. We also calculate the CLIP Score for each method shown in Tab. 1, which uses the CLIP [35] model to extract features and calculate cosine similarity between input text and generated image. Our method also achieves optimal result. In addition, our method also has a significant advantage in running time, as shown in Tab. 1. Our method is able to fine-tune model on the given text prompt in about 3 minutes, while other approaches require more time. Although 3DGAN-Inv costs comparable running time to ours, its portrait generation quality is poor and fails to generate freestyle portraits.
Figure 8: The ablation studies of our approach. (a) Ip2p generated results w/ or w/o inference optimization. (b) Different ablation experiments on the 3D Latent Feature Generator (3DLFG). (c) 3D shape outputs w/ or w/o dr loss.
@@ 4.4 Ablation study
Table 2: Quantitative ablation results on 3D latent feature generator.
@@
Ours 0.332
We conduct the ablation studies of our method from three aspects, which are shown in Fig. 8. First, when we do not optimize the Ip2p inference, some character prompts generate poor results, such as the example in (a) of Fig. 8). Second, we perform the ablation of the 3D latent feature generator in (b) of Fig. 8. 1) When we remove the 3D latent feature generator and directly fine-tuning the whole EG3D, we can achieve some stylization effect. However, the generation quality is somewhat poor, and this fine-tuning is inefficient, taking dozens of minutes. 2) When using our 3D latent feature generator without pre-training, training from scratch on the few-shot dataset will not be able to learn the correct 3D implicit representation. Because it is difficult to establish the mapping from the image to the triplane representation only using 3D inconsistent data. 3) When the $\mathbf{w}_{s}$ code of our 3D latent feature generator is a vector without spatial dimension, the learned 3D portrait is blurred and lacks details. Furthermore, we calculate the CLIP Score of different ablation models in Tab. 2, and our method achieves the best result. In addition, we ablate the density regularization (dr) loss used in model training, shown in (c) of Fig. 8. When the density regularization is not used, the synthesized portrait shape will have rough surfaces.
5 Limitation
@@
Our approach is based on two powerful pre-trained generative priors, which are the basis of our method’s ability to synthesize high-quality stylized 3D portraits. At the same time, our method is limited by both priors, especially Ip2p, which is unable to achieve perfect 3D-consistent portrait stylization in different viewpoints, so the final synthesized 3D portrait of our method differs slightly from the style generated by Ip2p. Meanwhile, some stylistic changes that differ significantly from the human portrait shape, such as ”Iron Man” and ”Stormtrooper” in Fig. 9, result in poorer quality 3D portraits compared to human portraits. Because when Ip2p generates this type of style, the stylization effect will be more different under different views.
@@ 6 Conclusions
This paper proposes a novel freestyle 3D-aware portrait synthesis framework based on compositional generative priors. We combine two pre-trained generative priors, EG3D and Ip2p, to quickly construct the out-of-distribution stylized multi-view portraits. We optimize Ip2p inference in order to stylize portraits at different views more freely and stable. We propose a 3D latent feature generator, which learns to map the style information from few-shot stylized portrait dataset to the 3D implicit representation. Equipping the generator with a pre-trained 3D renderer, we can generate 3D-consistent stylized portraits. A large number of high-quality 3D portrait synthesis results and comparison experiments with baselines show the superiority of our method.